C= R60 million


I= R280 million


Goverment= R310 million


C = 0.75



1. Calculate the Autonomous expenditure


2. Calculate the Equilibrium level of income.


3. Draw a graph to illustrate the Equilibrium

1
Expert's answer
2021-11-04T18:40:53-0400

1) AE = C+c(Y) +I+G

   = 60+0.75(Y) + 280+ 310

    = 650 + 0.75(Y) 

Here out of total expenditure , 650 is the autonomous expenditure. 


2) 

At equilibrium 

AE = Y 

60+0.75(Y) + 280+ 310 = Y 

0.25 Y = 650

Y = 2600

Equilibrium level of income is 2600.